WHAT IS THE SURFACE AREA OF A CUBE THAT IS 3 INCHES HIGH, 2 INCHES WIDE, AND 4 INCHES LONG?
Answer by checkley71(8403)

2(3*2)+4(2*4)=2*6+4*8=12+32=44in^2 answer.